# High-throughput assessment of defect-nuclear spin register controllability for quantum memory applications

### v1

@@ Abstract
Quantum memories play a key role in facilitating tasks within quantum networks and quantum information processing, including secure communications, advanced quantum sensing, and distributed quantum computing. Progress in characterizing large nuclear spin registers coupled to defect electronic spins has been significant, but selecting memory qubits remains challenging due to the multitude of possible assignments. Numerical simulations for evaluating entangling gate fidelities encounter obstacles, restricting research to small registers, while experimental investigations are time-consuming and often limited to well-understood samples. Here we present an efficient methodology for systematically assessing the controllability of defect systems coupled to nuclear spin registers. We showcase the approach by investigating the generation of entanglement links between defects in SiC and randomly selected sets of nuclear spins within the two-species (13C and 29Si) nuclear register. We quantify the performance of entangling gate operations and present the achievable gate fidelities, considering both the size of the register and the presence of unwanted nuclear spins. We find that some control sequences perform better than others depending on the number of target versus bath nuclei. This efficient approach is a guide for both experimental investigation and engineering, facilitating the high-throughput exploration of suitable defect systems for quantum memories.
@@ I Introduction
In this paper, we address this challenge by establishing an efficient method for assessing the controllability of the total system consisting of the defect spin coupled to an arbitrary multi-nuclear spin register. We do this by leveraging a recent formalism for characterizing multi-spin entanglement [28] that completely sidesteps the need to simulate many-body dynamics—a notoriously resource-intensive and time-consuming computational task. Our method enables one to rapidly and systematically identify suitable nuclear spins to use as memories, as well as optimal control sequences for manipulating them with high fidelity. While our method is general and can be applied to any defect, we showcase our approach with defects in silicon carbide coupled to two species of nuclear spins, 13C and 29Si. These defects are especially promising for networks due to the fact that they emit at or near telecom bands. We determine how electron-multi-nuclear gate fidelities depend on the number of target versus bath nuclei. We compare the performance of several leading DD sequences in this task, finding that the relative performance depends on the sizes of the nuclear spin memory register and the nuclear spin bath. We further show that the sign difference in the gyromagnetic ratios of the two nuclear spin species is beneficial for achieving high-fidelity multi-spin operations. Our results constitute an important step in comparing the utility of different defect systems by allowing memory controllability to be included systematically in such comparisons.
@@ IV Fidelity dependence on register and bath size
In this section, we quantify the effect of the register size and the spin bath on the gate fidelity in the case of a single silicon vacancy and a nuclear bath of 13C and 29Si nuclear spins. We begin by introducing a way to optimize the gate fidelity. We then implement this method for a monovacancy coupled to multiple nuclear spins, and, finally, we examine the case of a mixed bath where the nuclear spins have the same sign in gyromagnetic ratio.
@@ IV.1 Fidelity optimization
Figure 4: Gate infidelity $(1-F)$ versus the size of the bath (unwanted spins) and the size of the register (wanted or target spins) before (upper row) and after (bottom row) optimization over single-qubit rotations on the electronic spin. In (a) and (c) we plot the decimal logarithm of the statistical average of infidelity, i.e., $\log_{10}\left(\overline{1-F}\right)$, while in (b) and (d) we plot the variance of the infidelity on the same logarithmic scale $\log_{10}{\rm Var}\left(1-F\right)$. The numbers in each tile give the value of $\log_{10}(\overline{1-F})$ (left) and $\log_{10}{\rm Var}(1-F)$ (right). In the above simulations, we considered SiC samples with a silicon monovacancy with total spin $S=3/2$, and the electron qubit is defined by $(s_{0},s_{1})=(1/2,3/2)$. The HF values for the 13C and the 29Si were drawn from uniform distributions in the intervals $2\pi[10,200]\,\rm kHz$ and $2\pi[0.5,200]\,\rm kHz$, respectively. The entangling power threshold was set to $\epsilon_{p}^{\rm th}=0.85$, the magnetic field is $B=83\rm G$ leading to Larmor frequencies $(\omega_{L}^{C},\omega_{L}^{Si})=2\pi(88.8797,\,-70.2595)\,\rm kHz$, and the relative abundance was kept constant for all the cases, i.e., $\frac{{}^{29}\rm{Si}}{{}^{13}\rm{C}}=\frac{4.7\%}{1.1\%}=4.27$. Every data point on the heatmaps was obtained from 200 realizations, and for each trajectory we kept the combination $(\tau,N)$ that resulted in the lowest maximum unwanted one-tangle.
@@ IV.2 Average fidelity as a function of register and bath size
In the following simulations, we presume an electron spin with $S=3/2$ with the qubit defined using the $(s_{0},s_{1})=(1/2,3/2)$ spin states, which correspond to a silicon vacancy defect in SiC [9, 8, 7]. The nuclear Larmor frequencies are specified as $(\omega_{L}^{\rm C},\omega_{L}^{\rm Si})=2\pi(88.8797,\,-70.2595)\,\rm kHz$ for ${}^{13}\rm C$ and ${}^{29}\rm Si$ nuclei, respectively. In the absence of experimental data, we draw the HF values from uniform distributions in the range $2\pi[10,200]\,\rm kHz$ and $2\pi[0.5,200]\,\rm kHz$ for ${}^{13}\rm C$ and ${}^{29}\rm Si$ nuclei, respectively. These nuclei are weakly coupled since the HF parameters are smaller than $1/T_{2}^{*}$, which is typically a few hundred kHz for NV centers [42, 43] or, in general, when $A_{\|},A_{\bot}\ll 1\,\rm MHz$ [44] (approximately 1 MHz is also the electron linewidth for the neutral divacancy in SiC [10]). Moreover, to ensure the distinctness of spins chosen through random generation, we impose a condition on the discrepancy of HF values; we stipulate that at least one HF value ($A_{\|}$ or $A_{\perp}$) should deviate by at least $2\pi\cdot 10\,\rm kHz$ from the others. This condition is chosen to ensure the generation of sufficient nuclei within the HF range with distinct HF values in order to simulate a realistic silicon vacancy in a SiC sample. Our goal is to examine the dependence of the fidelity on several register-bath size combinations. To do so, we use the algorithm described in Sec. III.2, aiming to maximize the one-tangles of the target spins while minimizing those of the unwanted spins. Contrary to the simple example shown in Fig. 3, now will we explore the capabilities of five resonances ($k\in\{1,2,3,4,5\}$) of the CPMG, UDD3, and UDD4 pulse sequences.
@@ V Conclusions
Nuclear spins play a crucial role in spin-based solid-state platforms for quantum memories. To fully leverage their potential for creating large-scale quantum networks, a comprehensive investigation of the ability to control nuclear spin registers in all the promising candidate platforms should be conducted. Here, we developed a method for conducting such a high-throughput sorting of host materials and defects. As a concrete example to showcase our approach, we explored the performance of dynamical decoupling sequences such as CPMG, UDD3, and UDD4 in controlling the nuclear spins surrounding a defect in SiC. In each case, we quantified the expected performance in terms of the gate infidelity, and found an exponential deterioration not only with the bath size but also the register size. Additionally, we showed that some sequences perform better on average than others, depending on the register and bath size.

### v2

@@ Abstract
Quantum memories play a key role in facilitating tasks within quantum networks and quantum information processing, including secure communications, advanced quantum sensing, and distributed quantum computing. Progress in characterizing large nuclear spin registers coupled to defect electronic spins has been significant, but selecting memory qubits remains challenging due to the multitude of possible assignments. Numerical simulations for evaluating entangling gate fidelities encounter obstacles, restricting research to small registers, while experimental investigations are time-consuming and often limited to well-understood samples. Here we present an efficient methodology for systematically assessing the controllability of defect systems coupled to nuclear spin registers. We showcase the approach by investigating the generation of entanglement links between silicon monovacancy or divacancy centers in SiC and randomly selected sets of nuclear spins within the two-species (13C and 29Si) nuclear register. We quantify the performance of entangling gate operations and present the achievable gate fidelities, considering both the size of the register and the presence of unwanted nuclear spins. We find that some control sequences perform better than others depending on the number of target versus bath nuclei. This efficient approach is a guide for both experimental investigation and engineering, facilitating the high-throughput exploration of suitable defect systems for quantum memories.
@@ I Introduction
In this paper, we address this challenge by establishing an efficient method for assessing the controllability of the total system consisting of the defect spin coupled to an arbitrary multi-nuclear spin register. We do this by leveraging a recent formalism for characterizing multi-spin entanglement [28] that completely sidesteps the need to simulate many-body dynamics—a notoriously resource-intensive and time-consuming computational task. Our method enables one to rapidly and systematically identify suitable nuclear spins to use as memories, as well as optimal control sequences for manipulating them with high fidelity. While our method is general and can be applied to any defect, we showcase our approach with silicon monovacancy and divacancy defects in silicon carbide coupled to two species of nuclear spins, 13C and 29Si. These defects are especially promising for networks due to the fact that they emit at or near telecom bands. We determine how electron-multi-nuclear gate fidelities depend on the number of target versus bath nuclei. We compare the performance of several leading DD sequences in this task, finding that the relative performance depends on the sizes of the nuclear spin memory register and the nuclear spin bath. We further show that the sign difference in the gyromagnetic ratios of the two nuclear spin species is beneficial for achieving high-fidelity multi-spin operations. Our results constitute an important step in comparing the utility of different defect systems by allowing memory controllability to be included systematically in such comparisons.
@@ IV Fidelity dependence on register and bath size
In this section, we quantify the effect of the register size and the spin bath on the gate fidelity in the case of silicon monovacancies and divacancies coupled to a nuclear bath of 13C and 29Si nuclear spins. We begin by introducing a way to optimize the gate fidelity. We then implement this method for a monovacancy coupled to multiple nuclear spins, and, finally, we examine the case of a mixed bath where the nuclear spins have the same sign in gyromagnetic ratio.
@@ IV.1 Fidelity optimization
Figure 4: Gate infidelity $(1-F)$ versus the size of the bath (unwanted spins) and the size of the register (wanted or target spins) before (a,b) and after (c,d) optimization over single-qubit rotations on the electronic spin. In (a) and (c) we plot the decimal logarithm of the statistical average of infidelity, i.e., $\log_{10}\left(\overline{1-F}\right)$, while in (b) and (d) we plot the variance of the infidelity on the same logarithmic scale $\log_{10}{\rm Var}\left(1-F\right)$. The numbers in each tile give the value of $\log_{10}(\overline{1-F})$ (a,c) and $\log_{10}{\rm Var}(1-F)$ (b,d). In the above simulations, we considered SiC samples with a silicon monovacancy with total spin $S=3/2$, and the electron qubit is defined by $(s_{0},s_{1})=(1/2,3/2)$. The HF values for the 13C and the 29Si were drawn from uniform distributions in the intervals $2\pi[10,200]\,\rm kHz$ and $2\pi[0.5,200]\,\rm kHz$, respectively. The entangling power threshold was set to $\epsilon_{p}^{\rm th}=0.85$, the magnetic field is $B=83\rm G$ leading to Larmor frequencies $(\omega_{L}^{C},\omega_{L}^{Si})=2\pi(88.8797,\,-70.2595)\,\rm kHz$, and the relative abundance was kept constant for all the cases, i.e., $\frac{{}^{29}\rm{Si}}{{}^{13}\rm{C}}=\frac{4.7\%}{1.1\%}=4.27$. Every data point on the heatmaps was obtained from 200 realizations, and for each trajectory we kept the combination $(\tau,N)$ that resulted in the lowest maximum unwanted one-tangle.
@@ IV.2 Average fidelity as a function of register and bath size
Figure 6: Gate infidelity $(1-F)$ versus the size of the bath (unwanted spins) and the size of the register (wanted or target spins) upon optimization over single-qubit rotations on the electronic spin. In (a) we plot the decimal logarithm of the statistical average of infidelity, i.e., $\log_{10}\left(\overline{1-F}\right)$, while in (b) we plot the variance of the infidelity on the same logarithmic scale $\log_{10}{\rm Var}\left(1-F\right)$. The numbers in each tile give the value of $\log_{10}(\overline{1-F})$ and $\log_{10}{\rm Var}(1-F)$, respectively. In the above simulations, we considered SiC samples with a silicon divacancy with total spin $S=1$, and the electron qubit is defined by $(s_{0},s_{1})=(0,-1)$. The HF values for the 13C and 29Si nuclei were drawn from uniform distributions in the intervals $2\pi[10,200]\,\rm kHz$ and $2\pi[0.5,200]\,\rm kHz$, respectively. The entangling power threshold was set to $\epsilon_{p}^{\rm th}=0.85$, the magnetic field is $B=584\rm G$ leading to Larmor frequencies $(\omega_{L}^{C},\omega_{L}^{Si})=2\pi(625.37,\,-494.35)\,\rm kHz$, and the relative abundance was kept constant for all the cases, i.e., $\frac{{}^{29}\rm{Si}}{{}^{13}\rm{C}}=\frac{4.7\%}{1.1\%}=4.27$. Every data point on the heatmaps was obtained from 100 realizations, and for each trajectory we kept the combination $(\tau,N)$ that resulted in the lowest maximum unwanted one-tangle.
In the following simulations, we presume an electron spin with $S=3/2$ with the qubit defined using the $(s_{0},s_{1})=(1/2,3/2)$ spin states, which correspond to a silicon monovacancy defect in SiC [9, 8, 7]. The nuclear Larmor frequencies are specified as $(\omega_{L}^{\rm C},\omega_{L}^{\rm Si})=2\pi(88.8797,\,-70.2595)\,\rm kHz$ for ${}^{13}\rm C$ and ${}^{29}\rm Si$ nuclei, respectively. In the absence of experimental data, we draw the HF values from uniform distributions in the range $2\pi[10,200]\,\rm kHz$ and $2\pi[0.5,200]\,\rm kHz$ for ${}^{13}\rm C$ and ${}^{29}\rm Si$ nuclei, respectively. These nuclei are weakly coupled since the HF parameters are smaller than $1/T_{2}^{*}$, which is typically a few hundred kHz for NV centers [42, 43] or, in general, when $A_{\|},A_{\bot}\ll 1\,\rm MHz$ [44] (approximately 1 MHz is also the electron linewidth for the neutral divacancy in SiC [10]). Moreover, to ensure the distinctness of spins chosen through random generation, we impose a condition on the discrepancy of HF values; we stipulate that at least one HF value ($A_{\|}$ or $A_{\perp}$) should deviate by at least $2\pi\cdot 10\,\rm kHz$ from the others. This condition is chosen to ensure the generation of sufficient nuclei within the HF range with distinct HF values in order to simulate a realistic silicon vacancy in a SiC sample. Our goal is to examine the dependence of the fidelity on several register-bath size combinations. To do so, we use the algorithm described in Sec. III.2, aiming to maximize the one-tangles of the target spins while minimizing those of the unwanted spins. Contrary to the simple example shown in Fig. 3, now will we explore the capabilities of five resonances ($k\in\{1,2,3,4,5\}$) of the CPMG, UDD3, and UDD4 pulse sequences.
Next, we consider the case of a silicon divacancy defect in SiC with electron spin $S=1$ where the qubit is defined using the $(s_{0},s_{1})=(0,-1)$ spin states. Similarly to what we showed for the monovacancy center above, in Fig. 6 we show the statistical average of the infidelity and its variance over 100 realizations in a much stronger magnetic field, $B=584\,\rm G$, taken from Ref. [10]. Comparing Fig. 6(a) and Fig. 4(c), we observe that despite the difference in the spin states used to define the electronic spin qubit and the strength of the applied magnetic field, the fidelity deteriorates in the same fashion with the size of the register, while it decreases faster with the number of the unwanted spins. However, there are cases where the divacancy and a strong magnetic field lead to higher average fidelity; compare for instance the $N_{B}=1$ columns in Fig. 6(a) and Fig. 4(c). The last two features can be attributed to the fact that the combination of total spin $S=1$ and strong magnetic field lead to resonance times with lower dispersion, hence making it easier for DD pulses to create large registers but more difficult to decouple selected nuclear spins from a large bath. Also, low dispersion in resonance times can affect our ability to create a desired entangled state in the presence of a given number of unwanted spins (see Appendix D).
@@ V Conclusions
Nuclear spins play a crucial role in spin-based solid-state platforms for quantum memories. To fully leverage their potential for creating large-scale quantum networks, a comprehensive investigation of the ability to control nuclear spin registers in all the promising candidate platforms should be conducted. Here, we developed a method for conducting such a high-throughput sorting of host materials and defects. As a concrete example to showcase our approach, we explored the performance of dynamical decoupling sequences such as CPMG, UDD3, and UDD4 in controlling the nuclear spins surrounding a silicon monovacancy or divacancy defect in SiC. In each case, we quantified the expected performance in terms of the gate infidelity, and found an exponential deterioration not only with the bath size but also the register size. Additionally, we showed that some sequences perform better on average than others, depending on the register and bath size.
@@ Appendix C Gate fidelity optimization
Appendix D Success Probability
Here we present the total number of realizations we had to simulate to obtain the successful realizations used in the main text figures. As a successful realization, we define a realization in which we can find a DD sequence (CPMG or UDD) and a pair of pulse parameters $(\tau,N)$ such that a desired number of $N_{\rm R}$ nuclear spins have $\epsilon_{p}\geq\epsilon_{p}^{\rm th}$ and a desired number of $N_{\rm B}$ unwanted spins have $\epsilon_{p}<\epsilon_{p}^{\rm th}$. Any other realization is considered unsuccessful and we stop the simulation. For instance, we have an unsuccessful realization even in the case where the HF couplings are such that we cannot create a register of $N_{\rm R}$ spins surrounded by a bath of $N_{\rm B}$ spins, but we may be able to create a register of $N_{\rm R}-k$ spins surrounded by $N_{\rm B}+k$ spins (where $k$ is an integer). In Fig. 9 we show the total number of realizations needed for the silicon monovacancy, the divacancy, and the fictitious case of positive $\gamma_{{}^{29}\rm Si}$, where it becomes clear that the monovacancy in a low magnetic field has the highest success rate, while the fictitious bath leads to the lowest success rate. Furthermore, it becomes evident that the $(N_{\rm R},N_{\rm B})=(10,1)$ combination is the most challenging for all the cases, and especially for the fictitious one. However, except for the top left corner in the heatmaps, the success rate for all the other combinations is always equal to unity.
Figure 9: Total number of realizations simulated to obtain the results presented in the main text. Total number of realizations for (a) a silicon monovacancy defect in SiC, (b) a divacancy, and (c) a monovacancy in the fictitious nuclear spin bath where $\gamma_{{}^{29}\rm Si}>0$. The successful realizations for (a), (b), and (c) were 200, 100, and 100, respectively.